OpenStudy (anonymous):

How many atoms of gold are in 1 gram of gold? in order to start solving the question what should I look for 1st?

OpenStudy (anonymous):

well 1st you want to convert it to moles of gold. Do you know how to do this??

OpenStudy (chmvijay):

do you know atomic weight of gold its enough to find it

OpenStudy (anonymous):

i know i use the avogadros # 6.022x10^23 to convert to mole. but what do I divide by? just the one or the 197.0 mass?

OpenStudy (anonymous):

No you dont use avagadros # to convert to moles you use it to convert to convert to atoms

OpenStudy (anonymous):

you take 1 g of gold x 1 mol/molecular weight to get to mol of gold .. then you have 6.022 x 10^23 atoms per mole ... so you can multiple mols of gold times 6.022 x 10^23 atoms / 1 mol

OpenStudy (chmvijay):

197 gram of Gold has -6.023*23of arims then 1 gram of gold have =6.023*10^23 /197=

OpenStudy (chmvijay):

197 gram of Gold has -6.023*23of atoms then 1 gram of gold have =6.023*10^23 /197=

OpenStudy (chmvijay):

197 gram of Au =1 mole

OpenStudy (anonymous):

\[1 gram gold \times \frac{ 1 mol }{ molecular weight } \times \frac{ 6.022 \times 10^23 atoms }{ mol }\]

OpenStudy (chmvijay):

197 gram of Gold has -6.023*10^23of atoms then 1 gram of gold have =6.023*10^23 /197=

OpenStudy (anonymous):

Thank You now I get it.. thank you very much :)

OpenStudy (chmvijay):

ur welcome